If a 7.5 mL solution of HNO3 has a concentration of 0.05 M, what is its pH? Also, how
many grams of HCl are in the solution?
1
Expert's answer
2020-05-03T15:12:40-0400

HNO3 is a strong acid and dissociates completely or 100% .

ie., HNO3-------->H+ +NO3-

hence H+ ion concentration will be same as concentration of nitric acid

pH=-log[H+]=-log[5x10-2]=2-log5=2-0.7=1.3